Solve the equation. Give an exact solution.
4^x+6 = 7
This is a worksheet problem and I am getting stuck between using "ln" or "e" Your help would be appreciated.
Bonnie

Answer by mukhopadhyay(490)

4^x+6 = 7
=> 4^x = 1
=> x(log4) = log1 = 0 (by taking log on both sides)
=> x(log4) = 0
Since log4 > 0, x must be 0
Answer: x = 0